Two new computer codes are being developed to prevent unauthorized access to classified information. The first consists of six digits ( each chosen from 0 to 9); the second consists of three digits (from 0 to 9) followed by two letters ( A to Z), excluding I and O).
I) which code is better at preventing unauthorized access ( defined as breaking the code in one attempt)?
ii) if both codes are implemented, the first followed by the second, what is probability of gaining access in a single attempt?
1
Expert's answer
2021-09-08T18:30:21-0400

Number of combinations for the first code: "N_1=10^6=1000000."

Number of combinations for the second code: "N_2=10^3*24^2=576000."

i) Since "N_1>N_2", the first code is better.

ii) "N=N_1N_2=1000000*576000=576000000000=5.76*10^{11}."

"P=\frac{1}{N}=1.7361*10^{-12}."